### v1

# GPU-Accelerated Rule Evaluation and Evolution

Abstract.
This paper introduces an innovative approach to boost the efficiency and scalability of Evolutionary Rule-based machine Learning (ERL), a key technique in explainable AI. While traditional ERL systems can distribute processes across multiple CPUs, fitness evaluation of candidate rules is a bottleneck, especially with large datasets. The method proposed in this paper, AERL (Accelerated ERL) solves this problem in two ways. First, by adopting GPU-optimized rule sets through a tensorized representation within the PyTorch framework, AERL mitigates the bottleneck and accelerates fitness evaluation significantly. Second, AERL takes further advantage of the GPUs by fine-tuning the rule coefficients via back-propagation, thereby improving search space exploration. Experimental evidence confirms that AERL search is faster and more effective, thus empowering explainable artificial intelligence.
1. Introduction
Evolutionary Rule-based Machine Learning (ERL) systems harness evolutionary algorithms to optimize rule sets in many machine learning domains, including supervised learning where fitness is evaluated based on accuracy on a pre-existing dataset (Miikkulainen, 2023; Shahrzad et al., 2022). Graphics Processing Units (GPUs) have long been used to accelerate such algorithms, including Mixed Integer Non-Linear Programming (MINLP), Central Force Optimization (CFO), Genetic Algorithms (GA), and Particle Swarm Optimization (PSO), with speedups ranging from 7x to 10,000x (Singh et al., 2014). Further enhancing such optimistic performance claims (Jähne, 2016), this paper introduces a novel tensor-based representation of logical rules utilizing PyTorch. This representation not only enhances evaluation speed but also fosters local optimization capability via back-propagation on GPUs, thereby yielding significant performance improvements. The approach is validated empirically through preliminary experiments in the UCI breast cancer dataset, leading to a number of future work directions.
2. Background
Evolving transparent explainable rule sets based on extended propositional logic is a viable approach to creating explainable AI systems. Contrary to traditional black-box models, stakeholders can easily understand and audit such systems (Shahrzad et al., 2022; Hemberg et al., 2014). However, there is untapped potential in enhancing these processes through GPU acceleration. Given that evolutionary algorithms are computationally intensive, requiring rapid, iterative rule evaluation, GPU acceleration should be effective. Yet it is rarely used, presumably because rule-sets do not naturally map to GPUs the same way s black-box models such as neural networks.
| Features (only six included for simplicity): | ||
| mean radius, fractal dimension error, worst compactness, | ||
| worst symmetry, perimeter error, mean area | ||
| Rules: (Rule 1 has three conditions, Rule 2 has two) | ||
| Rule 1: torch.tensor([[1.0, 0, 0, 0, 0, 0], [0, 1.0, -0.38, 0, 0, 0], [0, 0, 0, 1.0, 0, 0]]), | ||
| Rule 2: torch.tensor([[0, 0, 0, 0, 1.0, 0], [0, 0, 0, 0, 0, 1.0]]) | ||
| Biases: | ||
| Rule 1: torch.tensor([-2.02, 0.0, -0.29]), # zero bias shows feature to feature comparison | ||
| Rule 2: torch.tensor([-0.43, -0.29]), |
() Numerical representation of the rule set before fine-tuning
() Rule set before fine-tuning
() The same rule set after fine-tuning
3. Method
The AERL approach transforms list-based propositional logic expressions (Shahrzad et al., 2022) into a computational model that can be efficiently processed using linear algebra and optimization techniques. By representing the rules as inequalities, AERL can process the logical conditions and their relationships in a structured numerical format. This conversion is pivotal because it allows the use of matrix multiplication to handle the evaluation of multiple rules simultaneously, thereby mapping the input features to the outcomes defined by the rules.
Specifically, each rule’s condition is translated into an inequality that, when applied to the input features through matrix multiplication, produces a vector (Figure 1). This vector represents the outcome of each rule for the given input, where the size of the vector corresponds to the number of classes or rules being evaluated. To interpret these outcomes in a binary fashion—true or false—AERL applies an activation function. This function clips negative values to zero, representing a false condition, and rounds positive values to one, representing a true condition. This binarization process is crucial for making discrete decisions based on the continuous outputs of matrix multiplication.
For the purpose of training or fine-tuning the model, a continuous activation function like the sigmoid can be employed. Unlike the clipping function used for inference, the sigmoid function provides a smooth gradient across all input values. This gradient is essential for gradient-based optimization techniques, allowing for the fine-tuning of the coefficients that scale the input features in the rules.
At each generation, parents are chosen for reproduction through tournament selection, and offspring generated with randomized coefficient values. Thus, back-propagation is only used to fine-tune the candidates’ fitness, i.e. it guides evolution through the Baldwin effect (Baldwin, 1896). Evolution takes place on the CPUs in all experiments; in the GPU experiments, the candidates are then loaded to the GPU for evaluation.
Figures 1 demonstrates this process. It shows a randomly generated rule set for the UCI breast cancer dataset, exhibiting an initial accuracy of 21.05%. Following fine-tuning via back-propagation over 50 epochs, the accuracy improved dramatically, exceeding 92.40%. The structure of the rules was preserved during the fine-tuning process, though the coefficients were adjusted.
4. Experimental Results
4.1. Experiments Setup
Experiments were conducted to evaluate the performance and efficiency of evolving PyTorch rules for the UCI Breast Cancer classification domain (Zwitter and Soklic, 1988). The key objective was to assess the impact of GPU acceleration and the application of fine-tuning through back-propagation on the evolutionary process. Experiments were run with and without fine-tuning, and on CPUs and GPUs, comparing performance and acceleration ratios. The evolution process was repeated across 10 separate runs to estimate statistical significance, with each run iterating through a fixed pool size of 100 candidate rule sets on Google Colab T4-GPU. In the fine-tuning experiments, 50 epochs of back-propagation was run on rule set coefficients.
4.2. Evolutionary Performance
The performance of the PyTorch rules was first measured in terms of accuracy improvements over generations. As shown in Figure 2, the application of fine-tuning significantly expedited the evolutionary progress, achieving statistically significant performance improvements (). This effect underscores the value of integrating gradient-based optimization methods within evolutionary algorithms to enhance the discovery of optimal rule sets.
4.3. Computational Efficiency
Figure 3 illustrates the acceleration ratios achieved through GPU computation. Notably, both with and without fine-tuning, the use of GPU resources consistently resulted in at least a fivefold speed increase over traditional CPU processing. This acceleration was observed despite variations in the number of rules and conditions per generation, highlighting the efficacy of GPU acceleration in expediting the evolutionary computation of PyTorch rules.
5. Discussion and Future Work
The efficacy and versatility of AERL can be enhanced in several ways. A primary focus will be on expanding the model to handle categorical data, a critical step for applicability across diverse datasets. Additionally, advanced features of rule-based methods such as time-lags and exponents can be incorporated to capture more complex patterns and temporal dynamics in the data. The Softmax activation function can be used to facilitate probabilistic classification, offering a nuanced understanding of class memberships (Hodjat et al., 2018). Tree-based rule evolution, akin to neural architecture search, is a potentially powerful extension, and allows leveraging more effective genetic operators (Qiu and Miikkulainen, 2023). These extensions should lead to significant advancements in the model’s ability to evolve complex rule sets, potentially leading to improvements in both explainability and accuracy.
6. Conclusion
The findings reinforce the premise that GPU-accelerated evolutionary computation, leveraged by fine-tuning techniques, can serve as a powerful tool in the development of explainable AI models.