### v2

# GPU-Accelerated Rule Evaluation and Evolution

Abstract.
This paper presents an extension to the EVOTER (Evolution of Transparent Explainable Rule Sets) framework, previously introduced to promote transparency and explainability in AI through the evolution of interpretable rule sets. While EVOTER demonstrated strong performance in generating compact and understandable rule-based models using a list-based grammar, the current work, Accelerated Evolutionary Rule-based Learning (AERL), builds upon this foundation by addressing scalability challenges. Specifically, AERL leverages GPU-accelerated computation and back-propagation fine-tuning within a PyTorch framework to significantly enhance the efficiency of rule evaluation and evolution. By introducing a tensorized numerical representation of rules and incorporating gradient-based optimization, AERL achieves superior computational speed and improved search-space exploration. Experimental results confirm the effectiveness of the approach in achieving faster convergence and better accuracy, offering a scalable path forward for transparent AI models.
1. Introduction
Evolutionary Rule-based Machine Learning (ERL) systems have been applied to a wide range of supervised learning tasks, where rule fitness is typically evaluated based on classification accuracy on a pre-existing dataset (Miikkulainen, 2023). A notable recent contribution in this space is EVOTER (Evolution of Transparent Explainable Rule Sets), which demonstrated how expressive rule sets can be evolved using a list-based grammatical framework (Shahrzad et al., 2024).
This paper presents an extension of EVOTER that addresses one of the primary bottlenecks in ERL systems: the computational cost of rule evaluation. Graphics Processing Units (GPUs) have long been used to accelerate optimization algorithms, including Mixed Integer Non-Linear Programming (MINLP), Central Force Optimization (CFO), Genetic Algorithms (GA), and Particle Swarm Optimization (PSO), with reported speedups ranging from 7x to 10,000x (Singh et al., 2014). Further enhancing such promising performance claims (Jähne, 2016), this work introduces a novel tensor-based representation of logical rules using the PyTorch framework. This representation enables massively parallel evaluation of rule sets on GPUs and supports local optimization through back-propagation, significantly accelerating both rule evaluation and search-space exploration.
By embedding rule sets into a numerical tensor format, the proposed method—Accelerated Evolutionary Rule-based Learning (AERL)—not only improves scalability but also leverages gradient-based fine-tuning to enhance evolutionary progress. Experimental validation on the UCI Breast Cancer dataset confirms the effectiveness of the approach, demonstrating substantial gains in both computational speed and model performance.
2. Background
Evolving transparent explainable rule sets based on extended propositional logic is a viable approach to creating explainable AI systems. Contrary to traditional black-box models, stakeholders can easily understand and audit such systems (Shahrzad et al., 2024; Hemberg et al., 2014). However, there is untapped potential in enhancing these processes through GPU acceleration. Given that evolutionary algorithms are computationally intensive, requiring rapid, iterative rule evaluation, GPU acceleration should be effective. Yet it is rarely used, presumably because rule-sets do not naturally map to GPUs the same way s black-box models such as neural networks.
| Features (only six included for simplicity): | ||
| mean radius, fractal dimension error, worst compactness, | ||
| worst symmetry, perimeter error, mean area | ||
| Rules: (Rule 1 has three conditions, Rule 2 has two) | ||
| Rule 1: torch.tensor([[1.0, 0, 0, 0, 0, 0], [0, 1.0, -0.38, 0, 0, 0], [0, 0, 0, 1.0, 0, 0]]), | ||
| Rule 2: torch.tensor([[0, 0, 0, 0, 1.0, 0], [0, 0, 0, 0, 0, 1.0]]) | ||
| Biases: | ||
| Rule 1: torch.tensor([-2.02, 0.0, -0.29]), # zero bias shows feature to feature comparison | ||
| Rule 2: torch.tensor([-0.43, -0.29]), |
() Numerical representation of the rule set before fine-tuning
() Rule set before fine-tuning
() The same rule set after fine-tuning
3. Method
The AERL approach transforms list-based propositional logic expressions (Shahrzad et al., 2024) into a computational model that can be efficiently processed using linear algebra and optimization techniques. By representing the rules as inequalities, AERL can process the logical conditions and their relationships in a structured numerical format. This conversion is pivotal because it allows the use of matrix multiplication to handle the evaluation of multiple rules simultaneously, thereby mapping the input features to the outcomes defined by the rules.
Specifically, each rule’s condition is translated into an inequality that, when applied to the input features through matrix multiplication, produces a vector (Figure 1). This vector represents the outcome of each rule for the given input, where the size of the vector corresponds to the number of classes or rules being evaluated. To interpret these outcomes in a binary fashion—true or false—AERL applies an activation function. This function clips negative values to zero, representing a false condition, and rounds positive values to one, representing a true condition. This binarization process is crucial for making discrete decisions based on the continuous outputs of matrix multiplication.
For the purpose of training or fine-tuning the model, a continuous activation function like the sigmoid can be employed. Unlike the clipping function used for inference, the sigmoid function provides a smooth gradient across all input values. This gradient is essential for gradient-based optimization techniques, allowing for the fine-tuning of the coefficients that scale the input features in the rules.
At each generation, parents are chosen for reproduction through tournament selection, and offspring generated with randomized coefficient values. Thus, back-propagation is only used to fine-tune the candidates’ fitness, i.e. it guides evolution through the Baldwin effect (Baldwin, 1896). Evolution takes place on the CPUs in all experiments; in the GPU experiments, the candidates are then loaded to the GPU for evaluation.
Figures 1 demonstrates this process. It shows a randomly generated rule set for the UCI breast cancer dataset, exhibiting an initial accuracy of 21.05%. Following fine-tuning via back-propagation over 50 epochs, the accuracy improved dramatically, exceeding 92.40%. The structure of the rules was preserved during the fine-tuning process, though the coefficients were adjusted.
4. Experimental Results
4.1. Experiments Setup
Experiments were conducted to evaluate the performance and efficiency of evolving PyTorch rules for the UCI Breast Cancer classification domain (Zwitter and Soklic, 1988). The key objective was to assess the impact of GPU acceleration and the application of fine-tuning through back-propagation on the evolutionary process. Experiments were run with and without fine-tuning, and on CPUs and GPUs, comparing performance and acceleration ratios. The evolution process was repeated across 10 separate runs to estimate statistical significance, with each run iterating through a fixed pool size of 100 candidate rule sets on Google Colab T4-GPU. In the fine-tuning experiments, 50 epochs of back-propagation was run on rule set coefficients.
4.2. Evolutionary Performance
The performance of the PyTorch rules was first measured in terms of accuracy improvements over generations. As shown in Figure 2, the application of fine-tuning significantly expedited the evolutionary progress, achieving statistically significant performance improvements (). This effect underscores the value of integrating gradient-based optimization methods within evolutionary algorithms to enhance the discovery of optimal rule sets.
4.3. Computational Efficiency
Figure 3 illustrates the acceleration ratios achieved through GPU computation. Notably, both with and without fine-tuning, the use of GPU resources consistently resulted in at least a fivefold speed increase over traditional CPU processing. This acceleration was observed despite variations in the number of rules and conditions per generation, highlighting the efficacy of GPU acceleration in expediting the evolutionary computation of PyTorch rules.
5. Discussion and Future Work
The efficacy and versatility of AERL can be enhanced in several ways. A primary focus will be on expanding the model to handle categorical data, a critical step for applicability across diverse datasets. Additionally, advanced features of rule-based methods such as time-lags and exponents can be incorporated to capture more complex patterns and temporal dynamics in the data. The Softmax activation function can be used to facilitate probabilistic classification, offering a nuanced understanding of class memberships (Hodjat et al., 2018). Tree-based rule evolution, akin to neural architecture search, is a potentially powerful extension, and allows leveraging more effective genetic operators (Qiu and Miikkulainen, 2023). These extensions should lead to significant advancements in the model’s ability to evolve complex rule sets, potentially leading to improvements in both explainability and accuracy.
6. Conclusion
The findings reinforce the premise that GPU-accelerated evolutionary computation, leveraged by fine-tuning techniques, can serve as a powerful tool in the development of explainable AI models.